# D-inverse constellations

### v1

@@ Abstract.
We give an algebraic characterisation of ordered groupoids, namely, we show that there is a categorical isomophism between the category of ordered groupoids and the category of $D$-inverse constellations. Here constellations are partial algebras in the sense that they possess a partial product, and a unary operation $D$. We consider constellations in which elements have a suitable notion of inverse, giving the notion of a D-inverse constellation.
@@ 1. Introduction
The Ehresmann-Schein-Nambooripad (ESN) Theorem establishes a categorical equivalence, indeed an isomorphism, between the category of inverse semigroups and the category of inductive groupoids; we refer the reader to [8] for an exposition of this result and its history. One can think of the ESN Theorem as replacing an algebraic structure by a small category equipped with an order that allows one to reconstruct the ‘lost’ algebraic information. Following the establishment of the ESN theorem, Lawson provided a correspondence along similar lines in [7] that connected so-called Ehresmann semigroups (bi-unary semigroups with both and domain and range-like operations) to certain types of ordered categories. Two-sided restriction semigroups are Ehresmann semigroups, and are covered as a special case in [7]. However, one-sided restriction semigroups are perhaps even more natural than two-sided ones since they provide the algebraic models for partial functions, so there was interest in obtaining an analogous result for them. The problem was the lack of a range operation, and more generally, the lack of sufficient left-right dual properties.
@@
In [6], the notion of a “true inverse” of an element in a left restriction semigroup $S$ was considered; $s\in S$ has true inverse $t$ if $st=D(s)$ and $ts=D(t)$. It was shown that true inverses are unique if they exist, and that if every element has a true inverse then $S$ is an inverse semigroup in which $s^{\prime}$ is the true inverse of $s$. Conversely, it is easy to show that every inverse semigroup gives rise to a left restriction semigroup in which every element has a true inverse, by setting $D(s)=ss^{\prime}$ for all $s\in S$ (where $s^{\prime}$ is the inverse of $s$). These constructions are easily seen to be mutually inverse, so one may view inverse semigroups as nothing but left restriction semigroups in which every element has a true inverse.
In the current work, we use this same idea to define analogues of ordered groupoids within the class of constellations. We show that these are nothing but ordered groupoids expressed in a different language.
In Section 2, we recall the required background notions. In Section 3, attention turns to the definitions and basic properties of D-regular and D-inverse constellations, defined by analogy with true inverses in left restriction semigroups. We show that small D-inverse constellations have a Cayley-style representation theorem in terms of one-to-one partial maps, generalising the Vagner-Preston theorem for inverse semigroups.
In Section 4, we show in Theorem 4.13 that D-inverse constellations are nothing but ordered groupoids “in disguise”. This raises the possibility of using D-inverse constellations as a tool to study ordered groupoids and hence invere semigroups. We do this by exploiting a known correspondence between constellations with range and ordered categories with restriction, established in [4].
In Section 5, we give an illustration of how one might use inverse constellations, by giving a construction of inverse constellations from semigroups with a distinguished set of idempotents. This construction generalises both Nambooripad’s construction of an ordered groupoid from a regular semigroup and Lawson’s construction of an ordered groupoid from the partial isometries of a semigroup with involution.
Inverse semigroups are not defined using a domain operation, although such an operation can be defined in any inverse semigroup, in terms of composition and inverse. So it is not surprising that inverse constellations may also be given a “domain-free” description, using the notion of a pre-constellation (these satisfy the defining laws of constellations that do not involve the domain operation). This is the subject of Section 6, where we show that there is a more general notion of “inverse pre-constellation” amongst which D-inverse constellations may be characterised in purely pre-constellation terms.
@@ 2.1. Notational conventions
Throughout, we generally write functions on the right of their arguments rather than the left, so “$xf$” rather than “$f(x)$”. Correspondingly, we write function compositions left to right, so that “$fg$” is “first $f$, then $g$”. An exception to this is unary operation application: if $D$ is a unary operation on the set $S$, we write $D(s)$ for $s\in S$ rather than $sD$.
@@ 2.2. Partial binary algebras
[…]
@@ 2.3. Categories and groupoids
[…]
@@
The identities $e,f$ in (Cat3) are easily seen to be unique, and we write $D(x)=e$ and $R(x)=f$. (Note that this is the opposite of the convention often used, and corresponds to the fact that we view a composition of functions $fg$ as “first $f$, then $g$”, discussed earlier.)
@@
In view of this, particularly when our category is small, we will often view a category as a partial algebra $\mathcal{C}=(C,\,\circ\,,D,R)$. With this viewpoint, a small category with a single identity (equivalently, all products defined) is simply a monoid.
A groupoid is a structure $(Q,\,\circ\,,\,^{\prime}\,,D,R)$ such that $(Q,\,\circ\,,D,R)$ is a category and is a unary operation on $Q$, such that for all $s\in Q$, $s\circ s^{\prime}=D(s)$ and $s^{\prime}\circ s=R(s)$. Again with this viewpoint, a small groupoid with a single identity is simply a group.
@@ 2.4. Ordered categories and groupoids
Following [7] (and using nomenclature consistent with that used there), we say a category $\mathcal{C}=(C,\,\circ,D,R)$ equipped with a binary relation $\leq$ on $C$ is an ordered category with restrictions if
@@
    if $x_{1}\leq x_{2}$ and $y_{1}\leq y_{2}$ and both $x_{1}\circ y_{1}$ and $x_{2}\circ y_{2}$ exist then $x_{1}\circ y_{1}\leq x_{2}\circ y_{2}$
@@
We remark that the term ‘ordered category’ may also mean a category satisfying the conditions above, together with the left-right dual of (OC8(i)).
@@
    $x\leq y$ implies $x^{\prime}\leq y^{\prime}$
@@
    Law (OC3) for ordered categories with restrictions holds
@@
    Law (OC8(i)) for ordered categories with restrictions holds.
@@ 2.5. Inductive groupoids and inverse semigroups
First, an inductive groupoid is an ordered groupoid $\mathcal{Q}$ in which $D(\mathcal{Q})$ forms a semilattice under the ordering. On the other hand, a semigroup $S$ is inverse if for every $a\in S$ there exists a unique $b\in S$ such that $a=aba$ and $b=bab$; usually, $b$ is denoted by $a^{\prime}$. In an inverse semigroup the idempotents of $S$ commute, that is, they form a semilattice (for the details, see [8]). The canonical example of an inverse semigroup is the symmetric inverse monoid $\mathscr{I}_{X}$ on a set $X$. The elements of $\mathscr{I}_{X}$ are partial bijections of $X$ and the composition $fg$ of $f$ and $g$ in $\mathscr{I}_{X}$ is that of partial functions.
@@ 2.6. Constellations
Since $e$ in (Const3) is unique given $x\in\mathcal{Q}$, we call it $D(x)$. It follows that $D(\mathcal{Q})=\{D(s)\mid s\in\mathcal{Q}\}=RI(\mathcal{Q}),$ the set of right identities of $\mathcal{Q}$. We call $D(\mathcal{Q})$ the projections of $\mathcal{Q}$. We adopt the usual convention of referring to the constellation $(\mathcal{Q},\cdot)$ simply as $\mathcal{Q}$ if there is no ambiguity. However, following the convention for categories, because $D$ may be viewed as a unary operation, we also often view constellations as partial algebras $(\mathcal{Q},\cdot\,,D)$. We say a constellation $\mathcal{Q}$ is small if $\mathcal{Q}$ is a set.
@@
We emphasise that a constellation $\mathcal{Q}$ is a (partial) algebra. However, from the algebraic structure we obtain an ordered structure. Specifically, we define the relation $s\leq t$ for $s,t\in\mathcal{Q}$ whenever $s=e\cdot t$ for some $e\in D(\mathcal{Q})$ (equivalently, $s=D(s)\cdot t$). The restriction to $D(\mathcal{Q})$, also denoted by $\leq$, is then given by $e\leq f$ if and only if $e\cdot f(=e)$ exists. A constellation $\mathcal{Q}$ is normal if for $\mbox{for all }e,f\in D(\mathcal{Q}),\mbox{ if $e\cdot f$ and $f\cdot e$ exist, then }e=f$. This is a property reminiscent of idempotents in a semigroup commuting, and relates to whether the given relations on $\mathcal{Q}$ are partial orders.
@@
Let $\mathcal{Q}$ be a constellation. Then the relation $\leq$ on $\mathcal{Q}$ is a quasiorder which we call the natural quasiorder; its restiction to $D(\mathcal{Q})$ is called the standard quasiorder. In both these cases, the quasiorder is a partial order if and only if $\mathcal{Q}$ is normal, and then we use “order” rather than “quasiorder”.
An important example of a small normal constellation, introduced in [1], is ${\mathcal{C}}_{X}$, consisting of partial functions on the set $X$, in which $s\cdot t$ is the usual composite of $s$ followed by $t$ provided $\operatorname{Im}(s)\subseteq\operatorname{Dom}(t)$, and undefined otherwise, and $D(s)$ is the restriction of the identity map on $X$ to $\operatorname{Dom}(s)$.
A subconstellation $Q$ of a constellation $\mathcal{Q}$ is a subset of $\mathcal{Q}$ that is closed under the constellation product wherever it is defined, and closed under $D$; $Q$ is then a constellation in its own right as is easily seen [3]. An important subconstellation of $\mathcal{C}_{X}$ consists of the elements which are one-one as (partial) maps; this yields the constellation $\mathcal{I}_{X}$, which has the same underlying set $I_{X}$ as $\mathscr{I}_{X}$, but a restricted set of products.
@@
A strong radiant $\rho:\mathcal{Q}_{1}\rightarrow\mathcal{Q}_{2}$ is a radiant for which, for all $s,t\in\mathcal{Q}_{1}$, $s\cdot t$ exists if and only if $s\rho\cdot t\rho$ does. We say $\rho:\mathcal{Q}_{1}\rightarrow\mathcal{Q}_{2}$ is an isomorphism if $\rho$ is a strong bijective radiant, and an embedding if it is a strong injective radiant. If $\rho:\mathcal{Q}_{1}\rightarrow\mathcal{Q}_{2}$ is an embedding then the image of $\rho$ is a subconstellation of $\mathcal{Q}_{2}$ that is isomorphic to $\mathcal{Q}_{1}$.
@@ 3.1. D-regularity
As a step towards defining D-inverse constellations, we have the following, motivated by corresponding notions of regularity in semigroup and ring theory. Specifically, a semigroup $S$ is regular if for all $a\in S$ there exists $b\in S$ such that $a=aba$; the element $ab$ is then an idempotent and a left identity for $a$. Recall that a semigroup is inverse if and only if it is regular and its idempotents form a semilattice [8].
@@
In a D-regular constellation $\mathcal{P}$, we have that $E(\mathcal{P})\subseteq RI(\mathcal{P})$ and so the two sets are equal (and so also to $D(\mathcal{P})$).
@@
For $e\in E(P)$, suppose $e^{\prime}$ is such that $e\cdot e^{\prime}=D(e)$. Then $D(e)=e\cdot e^{\prime}=(e\cdot e)\cdot e^{\prime}=e\cdot(e\cdot e^{\prime})=e\cdot D(e)=e$, so $e\in D(P)$. ∎
Let $\mathcal{P}$ be a constellation, let $a\in P$ be $D$-regular, and let $b\in P$ be such that $a\cdot b=D(a)$. Then as $D(a)\cdot a$ exists we certainly have $a=(a\cdot b)\cdot a$, but we need not have that $b\cdot a$ exists. In the case where it does, then $b\cdot a\in E(P)$ and also $b=(b\cdot a)\cdot b$. If further $\mathcal{P}$ is regular, then by Proposition 3.2 we have that $b\cdot a=D(b)$. The following example, showing that in the above scenario we need not have $b\cdot a\in D(P)$, is constructed from partial functions in a a typical manner.
Example 3.3.
Let $X=\{x,y,z\}$ and let $\mathcal{P}$ be the subconstellation of $\mathcal{C}_{X}$ with elements:
$a=\{(x,y)\},b=\{(y,x),(z,z)\},e=\{(x,x)\}\mbox{ and }f=\{(y,y),(z,z)\}.$
It is clear that $a$ is D-regular and $a\cdot b=D(a)=e$, but $b\cdot a$ does not exist and $b$ is not D-regular.
@@
The pre-constellation $\mathcal{P}$ is right cancellative if for all $a,b,c\in P$, if $a\cdot c=b\cdot c$, then $a=b$.
@@
Suppose $\mathcal{Q}=(Q,\,\cdot\,,D)$ is a D-regular constellation. If $a,b,c\in Q$ and $a\cdot c=b\cdot c$ then since there exists $d\in Q$ for which $c\cdot d=D(c)$, both $a\cdot(c\cdot d)$ and $b\cdot(c\cdot d)$ exist. Then $a\cdot(c\cdot d)=(a\cdot c)\cdot d=(b\cdot c)\cdot d=b\cdot(c\cdot d)$, so $a\cdot D(c)=b\cdot D(c)$, giving that $a=b$. ∎
Let $X$ be a set. The subconstellation of one-to-one partial functions $\mathcal{I}_{X}=(I_{X},\,\cdot\,,D)$ of $\mathcal{C}_{X}$ is right cancellative and normal (recalling that $s\cdot t$ is interpreted as “first $s$ then $t$”). Conversely, there is a “Cayley theorem” for right cancellative constellations, as follows.
@@
Suppose $\mathcal{P},\mathcal{Q}$ are constellations and $\mathcal{P}$ is D-regular, with $\rho:\mathcal{P}\rightarrow\mathcal{Q}$ a surjective radiant. Then $\mathcal{Q}$ is D-regular.
@@
Suppose $s\cdot t=D(s)$, where $s,t\in P$. Then $D(s\rho)=D(s)\rho=(s\cdot t)\rho=s\rho\cdot t\rho$, so, since $\rho$ is surjective, we deduce that $\mathcal{Q}$ is D-regular. ∎

### v2

@@ Abstract.
Constellations are partial algebras in the sense that they possess a partial product, and a unary operation modelling domain. They were first used to give an ESN-style theorem for left restriction semigroups in terms of so-called inductive constellations. Here, we consider constellations in which elements have a suitable notion of inverse, giving the notion of a D-inverse constellation. We show that there is a categorical isomorphism between the category of ordered groupoids and the category of D-inverse constellations. This may be viewed as a generalisation of the ESN theorem, which relates the category of inductive groupoids to the category of inverse semigroups.
@@ 1. Introduction
The Ehresmann-Schein-Nambooripad (ESN) Theorem establishes a categorical equivalence, indeed an isomorphism, between the category of inverse semigroups and the category of inductive groupoids; we refer the reader to [9] for an exposition of this result and its history. One can think of the ESN Theorem as replacing an algebraic structure by an ordered one, which in this case is a small category equipped with an order that allows one to reconstruct the ‘lost’ algebraic information. Following the establishment of the ESN theorem, Lawson provided a correspondence along similar lines in [8] that connected so-called Ehresmann semigroups (bi-unary semigroups with both and domain and range-like operations) to certain types of ordered categories. Two-sided restriction semigroups are Ehresmann semigroups, and are covered as a special case in [8]. However, one-sided restriction semigroups are perhaps even more natural than two-sided ones since they provide the algebraic models for partial functions, so there was interest in obtaining an analogous result for them. The problem was the lack of a range operation, and more generally, the lack of sufficient left-right dual properties.
@@
In [7], the notion of a “true inverse” of an element in a left restriction semigroup $S$ was considered: $s\in S$ has true inverse $t$ if $st=D(s)$ and $ts=D(t)$. It was shown that true inverses are unique if they exist, and that if every element has a true inverse then $S$ is an inverse semigroup in which $s^{\prime}$ (that is, the inverse of $s$ in the inverse semigroup) is the true inverse of $s$. Conversely, it is easy to show that every inverse semigroup gives rise to a left restriction semigroup in which every element has a true inverse, by setting $D(s)=ss^{\prime}$ for all $s\in S$. These constructions are easily seen to be mutually inverse, so one may view inverse semigroups as nothing but left restriction semigroups in which every element has a true inverse.
Our main result, Theorem 4.14, uses an analogous notion of D-inverse to define analogues of ordered groupoids within the class of constellations. This reframing of an ordered structure to a purely algebraic one gives a new viewpoint, and allows alternative approaches to the theory of ordered groups and of inverse semigroups, already demonstrated in [6].
In Section 2, we recall the required background notions. In Section 3, attention turns to the definitions and basic properties of D-regular and D-inverse constellations, defined by analogy with left restriction semigroups in which each element has a true inverse. We show that small D-inverse constellations have a Cayley-style representation theorem in terms of one-to-one partial maps, generalising the Vagner-Preston theorem for inverse semigroups.
In Section 4, we show in Theorem 4.14 that D-inverse constellations are nothing but ordered groupoids “in disguise”. We do this by exploiting a known correspondence between constellations with range and ordered categories with restriction, established in [4]. This correspondence raises the possibility of using D-inverse constellations as a tool to study ordered groupoids and hence invere semigroups.
Indeed, in Section 5, we give an illustration of how one might use D-inverse constellations, by giving a construction of D-inverse constellations from semigroups with a distinguished set of idempotents. This construction generalises both Nambooripad’s construction of an ordered groupoid from a regular semigroup and Lawson’s construction of an ordered groupoid from the partial isometries of a semigroup with involution.
Inverse semigroups are not defined using a domain operation, although such an operation can be defined in any inverse semigroup, in terms of composition and inverse. So it is not surprising that D-inverse constellations may also be given a “domain-free” description, using the notion of a pre-constellation (these satisfy the defining laws of constellations that do not involve the domain operation). This is the subject of Section 6, where we show that there is a more general notion of “inverse pre-constellation” amongst which D-inverse constellations may be characterised in purely pre-constellation terms.
@@ 2. Background
We set out the necessary notions for this article, and clarify our notational conventions.
@@ 2.1. Notational conventions
Throughout, we generally write functions on the right of their arguments rather than the left, so “xfxf” rather than “f(x)f(x)”. Correspondingly, we write function compositions left to right, so that “fgfg” is “first $f$, then gg”. An exception to this is unary operation application: if $D$ is a unary operation on the set $S$, we write $D(s)$ for $s\in S$.
@@ 2.2. Partial binary algebras
Let $\mathcal{C}=(C,\,\cdot\,)$ be a class $C$ ((often a set)) equipped with a partial binary operation $\cdot$.
@@ 2.3. Categories and groupoids
For the purposes of this paper it is convenient to give an object-free formulation of a category, as occurs in the other works in this area, for example [3] and [8], which have brought us to this point in the research development. Recall that in this framework a category $\mathcal{C}=(C,\,\circ\,)$ is a class $C$ equipped with a partial binary operation $\circ$ satisfying the following:
@@
The identities $e,f$ in (Cat3) are easily seen to be unique, and we write $D(x)=e$ and $R(x)=f$. (Note that this is the opposite of the convention often used, and corresponds to the fact that we view a composition of functions $fg$ as “first $f$, then gg”, discussed earlier.)
@@
A functor $\rho$ from $\mathcal{C}$ to $\mathcal{D}$, where $\mathcal{C}=(C,\,\circ\,)$ and $\mathcal{D}=(D,\,\circ\,)$ are categories, is a map $\rho:C\rightarrow D$ such that for all $x\in\mathcal{C}$, $D(x)\rho=D(x\rho)$, $R(x)\rho=R(x\rho)$, and if $x\circ y$ exists in $\mathcal{C}$, then it equals $x\rho\circ y\rho$ (which must exist); that this is equivalent to the usual object-morphism notion of functor follows from the definitions.
In view of this, particularly when our category is small, we will often view a category as a partial algebra $\mathcal{C}=(C,\,\circ\,,D,R)$. With this viewpoint, a small category with a single identity (equivalently, in which all products are defined) is nothing but a monoid.
[…]
@@ 2.4. Ordered categories and groupoids
Following [8] (and using nomenclature consistent with that used there), we say a category $\mathcal{C}=(C,\,\circ,D,R)$ equipped with a binary relation $\leq$ on $C$ is an ordered category with restrictions if:
@@
    if $x_{1}\leq x_{2}$ and $y_{1}\leq y_{2}$ and both $x_{1}\circ y_{1}$ and $x_{2}\circ y_{2}$ exist, then $x_{1}\circ y_{1}\leq x_{2}\circ y_{2}$;
@@
(We remark that the term ‘ordered category’ is used in the literature to mean a category satisfying the conditions above, together with the left-right dual of (OC8(i)).)
@@
    $x\leq y$ implies $x^{\prime}\leq y^{\prime}$;
@@
    Law (OC3) for ordered categories with restrictions:
@@
    Law (OC8(i)) for ordered categories with restrictions.
@@ 2.5. Inductive groupoids and inverse semigroups
First, an inductive groupoid is an ordered groupoid $\mathcal{Q}$ in which $D(\mathcal{Q})$ forms a semilattice under the ordering. On the other hand, a semigroup $S$ is inverse if for every $a\in S$ there exists a unique $b\in S$ such that $a=aba$ and $b=bab$; usually, $b$ is denoted by $a^{\prime}$. In an inverse semigroup the idempotents of $S$ commute, and thus form a semilattice under the ordering $a\leq b$ if $ab=a$, where meets are given by the semigroup product. (For the details, see [9].) The canonical example of an inverse semigroup is the symmetric inverse monoid $\mathscr{I}_{X}$ on a set $X$. The elements of $\mathscr{I}_{X}$ are partial bijections of $X$ and the composition $fg$ of $f$ and $g$ in $\mathscr{I}_{X}$ is that of partial functions.
@@ 2.6. Constellations
Since $e$ in (Const3) is unique given $x\in\mathcal{Q}$, we call it $D(x)$. It follows that $D(\mathcal{Q})=\{D(s)\mid s\in\mathcal{Q}\}=RI(\mathcal{Q}),$ the set of right identities of $\mathcal{Q}$. We call $D(\mathcal{Q})$ the projections of $\mathcal{Q}$. We adopt the usual convention of referring to the constellation $(\mathcal{Q},\cdot)$ simply as $\mathcal{Q}$ if there is no ambiguity. However, following the convention for categories, and since $D$ may be viewed as a unary operation, we also often view constellations as partial algebras $(\mathcal{Q},\cdot\,,D)$. We say a constellation $\mathcal{Q}$ is small if $\mathcal{Q}$ is a set.
@@
[…]
@@
Let $\mathcal{Q}$ be a constellation. Then the relation $\leq$ on $\mathcal{Q}$ is a quasiorder which we call the natural quasiorder; its restriction to $D(\mathcal{Q})$ is called the standard quasiorder. In both these cases, the quasiorder is a partial order if and only if $\mathcal{Q}$ is normal, and then we use “order” rather than “quasiorder”.
Certainly not all constellations are normal. Indeed, any quasi-ordered set $(P,\leq)$ carries the structure of a constellation $\mathcal{P}=(P,\cdot)$, where $x\cdot y$ is defined if and only if $x\leq y$ and then $x\cdot y=x$ [1]; this constellation is normal if and only if $(P,\leq)$ is a poset. An important example of a small normal constellation, introduced in [1], is ${\mathcal{C}}_{X}$, with underlying set consisting of the partial functions on the set $X$, in which $s\cdot t$ is the usual composite of $s$ followed by $t$ provided $\operatorname{Im}(s)\subseteq\operatorname{Dom}(t)$, and undefined otherwise, and $D(s)$ is the restriction of the identity map on $X$ to $\operatorname{Dom}(s)$.
A subconstellation $\mathcal{P}$ of a constellation $\mathcal{Q}$ is a subset of $\mathcal{Q}$ that is closed under the constellation product wherever it is defined, and closed under $D$; then $\mathcal{P}$ is a constellation in its own right as is easily seen [3]. An important subconstellation of $\mathcal{C}_{X}$ consists of the elements which are one-one as (partial) maps; this yields the constellation $\mathcal{I}_{X}$, which has the same underlying set $I_{X}$ as $\mathscr{I}_{X}$, but a restricted set of products.
@@
A strong radiant $\rho:\mathcal{Q}_{1}\rightarrow\mathcal{Q}_{2}$ is a radiant for which, for all $\in\mathcal{Q}_{1}, sts\cdot t$ exists if and only if $s\rho\cdot t\rho$ does. We say $\rho:\mathcal{Q}_{1}\rightarrow\mathcal{Q}_{2}$ is an isomorphism if $\rho$ is a strong bijective radiant, and an embedding if it is a strong injective radiant. Note that an isomorphism is precisely a radiant with a radiant inverse. If $\rho:\mathcal{Q}_{1}\rightarrow\mathcal{Q}_{2}$ is an embedding then the image of $\rho$ is a subconstellation of $\mathcal{Q}_{2}$ that is isomorphic to $\mathcal{Q}_{1}$.
@@ 3. D-regular and D-inverse constellations
In this section we introduce the central new notion of this article, namely that of a D-inverse constellation.
@@ 3.1. D-regularity
Recall that a semigroup $S$ is regular if for all $a\in S$ there exists $b\in S$ such that $a=aba$; the element $ab$ is then an idempotent and a left identity for $a$. If a regular semigroup is inverse then clearly it is regular and we have pointed out that its idempotents commute; in fact the converse is also true [9]. We here consider a suitable variant of this notion for constellations.
@@
In a D-regular constellation $\mathcal{P}$, we have that $E(\mathcal{P})\subseteq RI(\mathcal{P})$ and so the two sets are equal ((and hence both equal $D(\mathcal{P})$)).
@@
Pick $e\in E(P)$, and suppose that $e^{\prime}$ is such that $e\cdot e^{\prime}=D(e)$. Then $D(e)=e\cdot e^{\prime}=(e\cdot e)\cdot e^{\prime}=e\cdot(e\cdot e^{\prime})=e\cdot D(e)=e$, so $e\in D(P)$. ∎
@@
The constellation $\mathcal{P}$ is right cancellative if $a\cdot c=b\cdot c$ implies that $a=b$, for any $a,b,c\in P$.
@@
Suppose that $\mathcal{Q}=(Q,\,\cdot\,,D)$ is a D-regular constellation. If $a,b,c\in Q$ and $a\cdot c=b\cdot c$ then since there exists $d\in Q$ for which $c\cdot d=D(c)$, both $a\cdot(c\cdot d)$ and $b\cdot(c\cdot d)$ exist. Then $a\cdot(c\cdot d)=(a\cdot c)\cdot d=(b\cdot c)\cdot d=b\cdot(c\cdot d)$, so $a\cdot D(c)=b\cdot D(c)$, giving that $a=b$. ∎
Let $X$ be a set. The subconstellation of one-to-one partial functions $\mathcal{I}_{X}=(I_{X},\,\cdot\,,D)$ of $\mathcal{C}_{X}$ is right cancellative and normal (recalling that $s\cdot t$ is interpreted as “first $s$ then tt”). Conversely, there is a “Cayley theorem” for right cancellative constellations, as follows.
@@
Suppose that $\mathcal{P},\mathcal{Q}$ are constellations and $\mathcal{P}$ is D-regular, and that $\rho:\mathcal{P}\rightarrow\mathcal{Q}$ is a surjective radiant. Then $\mathcal{Q}$ is D-regular.
@@
Suppose that $s\cdot t=D(s)$, where $s,t\in P$. Then $D(s\rho)=D(s)\rho=(s\cdot t)\rho=s\rho\cdot t\rho$, so, since $\rho$ is surjective, we deduce that $\mathcal{Q}$ is D-regular. ∎